# Diffusion Models for Computational Design at the Example of Floor Plans

### v1

@@ 5 Conclusion
In this paper we evaluate the use of diffusion models for computational design at the example of floor plans. We analyse the current state of the art in the field and explain diffusion model approaches. We identify the main limitations and develop a novel approach to improve fine-tuning of diffusion models trough semantic encoding. We analyse the workflows for using the diffusion models qualitatively and evaluate the performance of different diffusion models encodings quantitatively in a large set of experiments.
Our experiments show that current diffusion models already create floor plans and that results can be improved significantly with fine-tuning. Particularly our proposed semantic encoding improved validity of the generated floor plans to 90 %. However, our experiments also unveiled several shortcomings in the current diffusion models of which many track back to the lack of semantic understanding. Therefore, we identified several future directions for research in this new field of computation design that may in future automate many detailed planning steps. It is unclear yet, if we can get there with image-based diffusion models or if we need to start working on BIM-based Diffusion Models to unlock their semantic potential.
@@ 6 Declarations
The code for generating the training samples as as well as all validation queries will be open sourced on acceptance of the paper.
Funding
No funding was received for conducting this study.

### v2

@@ 8 Conclusion
In this paper we evaluated the use of diffusion models for computational design for the example of floor plans. We analysed the current state of the art in the field and explained diffusion model approaches. We identified the main limitations and developed a novel approach to improve fine-tuning of diffusion models through semantic encoding. We analysed the workflows for using the diffusion models qualitatively and evaluate the performance of different diffusion model encodings quantitatively in a large set of experiments. Architectural quality of the floor plans was not evaluated for now, and would need a separate study, for example involving expert reviews.
Our experiments show that current diffusion models can already create floor plans and that results can be improved significantly with fine-tuning. Particularly our proposed semantic encoding improved validity of the generated floor plans to 90 %. However, our experiments also unveiled several shortcomings in the current diffusion models, of which many track back to the lack of semantic understanding. Therefore, we identified several future directions for research in this new field of AI-driven computational design, which may in future ultimately be able to automate many detailed planning steps. It is as of yet unclear whether we can get there with just this is possible to achieve through appropriately powerful image-based diffusion models, or whether we need to start working on BIM-based Diffusion Models to unlock their semantic potential.
@@ 9 Declarations
The code for generating the training samples as as well as all validation queries are available on https://github.com/AI4SC/bim-diffusion-models.